# SONICS: Synthetic Or Not - Identifying Counterfeit Songs

### v2

@@ 4.3.2 Human-AI Benchmark
To evaluate Human performance in comparison to AI-based models, we selected a subset of 520 samples from our large test data. This evaluation employed a dynamic scoring system, similar to LMSYS [4], allowing public participation and live leaderboard updates, which will be made publicly available upon acceptance of this paper. Three human participants were involved in this benchmark, with their performance summarized in Table 4. In contrast to the AI benchmark using short (5 sec) or long (120 sec) audio samples, this human benchmark employed 25 sec clips. This choice stems from the observation that short clips hinder human identification due to subtle inaudible artifacts easily detected by AI, while longer clips don’t necessarily improve human performance due to the difficulty in leveraging long-range temporal dependencies.
As shown in Table 4, AI-based methods consistently outperform human participants across all test partitions. However, both humans and AI models struggled most with Udio algorithms, particularly Udio-32, where human sensitivity dropped to 23.4%. Conversely, Suno algorithms, especially Suno-v3.5, were easier to detect, with a human sensitivity of 81.7%. This mirrors the findings in the AI benchmark, where models demonstrated higher specificity than sensitivity, indicating greater ease in identifying real songs compared to fake ones. Further analysis revealed distinct characteristics within real and fake song audio patterns. For instance, Suno algorithms often produced synthetic or mechanical-sounding vocals, while Udio-32 algorithm occasionally created the "Karaoke effect." Furthermore, Udio algorithms demonstrated the ability to create songs with multiple voices and higher notes, a feature absent in Suno algorithms. On the other hand, real songs exhibited unique elements like a wide range of notes, diverse timbre, complex rhythms, and clear vocals with accented pronunciation. They also featured distinctive sounds like flutes and finger snaps, which are challenging for fake songs to replicate.
@@
Algorithm (Sens.) Suno v2 0.65 0.80 0.72 0.64 0.54 0.69
Suno v3 0.99 0.96 0.99 0.98 0.98 0.75
Suno v3.5 0.99 0.98 0.99 0.99 0.99 0.82
Udio 32 0.67 0.56 0.33 0.23 0.18 0.23
Udio 130 0.99 0.87 0.99 0.99 0.99 0.55
Fake Type (Sens.) Half Fake 0.91 0.93 0.92 0.90 0.88 0.71
Mostly Fake 0.84 0.79 0.72 0.69 0.64 0.66
Full Fake 0.83 0.77 0.72 0.68 0.64 0.63
Real (Spec.) 0.98 0.80 0.94 0.95 0.97 0.78
Fake (Sens.) 0.85 0.82 0.75 0.72 0.66 0.66
Overall (F1) 0.92 0.82 0.83 0.80 0.78 0.71
Table 4: Human-AI Benchmark
4.3.3 Efficiency Benchmark
To comprehensively evaluate the efficiency of the proposed SpecTTTra model alongside other methods, we measure various metrics across different song lengths using a P100 16GB GPU. The metrics considered include Speed (A/S $\rightarrow$ Audio per Second), Floating Point Operations (FLOPs), GPU Memory Consumption (Mem.) during the forward pass with a batch size of 14, activation count (# Act.), and parameter count (# Param.). The results are summarized in Table 5. Our analysis reveals that while ViT is the fastest model for 5-sec songs, it becomes the slowest for 120-sec songs and exhibits significant memory consumption, rendering it impractical for longer sequences. However, ViT remains the most efficient in terms of parameter count across both short and long songs. On the other hand, ConvNeXt, despite its strong detection performance, is the most resource-intensive model in terms of memory usage, FLOPs, and parameter count. In contrast, the proposed SpecTTTra model variants demonstrate remarkable efficiency, particularly in longer songs. Although they may not excel in short audio segments, the SpecTTTra-$\gamma$ variant is nearly 3 times faster than ViT and 2 times faster than ConvNeXt for 120-sec songs. Moreover, it consumes memory 3 times less than ViT and 6 times less than ConvNeXt, highlighting its superior efficiency. Thus, while SpecTTTra secures the 2nd position in the detection benchmark, it offers the best trade-off between detection and efficiency performance, making it the most balanced model overall.
@@ 4.3.3 Efficiency Benchmark
5 ConvNeXt 141 1.5 0.5 4 28
ViT 168 1.1 0.2 2 17
SpecTTTra-$\gamma$ 164 0.7 0.1 1 17
SpecTTTra-$\beta$ 162 1.1 0.2 2 17
SpecTTTra-$\alpha$ 153 2.8 0.6 6 17
120 ConvNeXt 39 43.1 13.6 129 28
ViT 34 31.7 6.2 67 17
SpecTTTra-$\gamma$ 97 10.1 1.9 20 24
SpecTTTra-$\beta$ 80 13.9 2.6 29 21
SpecTTTra-$\alpha$ 48 23.6 4.6 50 19
Table 5: Efficiency Benchmark
5 Limitations and Future Work
The real songs in the proposed SONICS dataset are queried from YouTube dynamically using their title and artist name, which introduces the possibility of not retrieving the correct audio. Through manual analysis of random samples, we estimate that this issue affects approximately 0.5% of the dataset. Also, due to the copyright of the real songs, we are unable to make them public. However, our dataset contains YouTube links for each song, which can be used to download the song easily using API [2]. Another limitation arises from the fake songs generated by the Udio platform, which cannot include lyrics from real songs. This restricts comprehensive evaluation in the Half Fake songs category, as it only contains fake songs from the Suno platform. Furthermore, the dataset currently includes only English-language songs, which may restrict the scope of our evaluation. In future work, we plan to extend our dataset to multiple languages to enable a more comprehensive analysis. Additionally, our current benchmarks are based solely on Mel Spectrogram input. Moving forward, we aim to incorporate raw audio and explore other feature extraction methods, such as LFCC and MFCC, to enhance the robustness of our evaluations.
@@ 6 Conclusion
In this paper, we introduced SONICS, a comprehensive dataset for end-to-end synthetic song detection, addressing limitations in existing datasets, such as lack of music diversity, short duration, and most importantly, the absence of end-to-end AI-generated songs. Moreover, we proposed the SpecTTTra model, which efficiently captures long-range temporal relationships in songs, achieving competitive performance to existing popular models while reducing computational costs significantly. Through extensive experiments, we established both AI-based and human benchmarks, demonstrating the dataset’s effectiveness in advancing synthetic song detection research. Our work lays the foundation for improved methods for ensuring fair attribution and protecting the integrity of musical artistry.
7 Acknowledgment
We express our gratitude to the Suno and Udio platforms, which we used for generating end-to-end synthetic songs. We also appreciate the GPU credits provided by Jarvislabs.ai, which were partially used for training our model.
@@ 7 Acknowledgment
Appendix A Dataset Statistics
@@
Figure 5: Long Context Correlation
Validation Data: In Table 6, the distribution of the labels in validation data can be seen. For proper evaluation, there is are 104 Full Fake, 235 Half Fake and 1,934 Mostly Fake songs, cumulating into 2,273 Fake songs. In comparison, there are 2,167 Real songs taken from YouTube in the validation data.
Table 6: Valid Data Summary
Label Source Count
Full Fake Suno-v2 16
Suno-v3 37
Suno-v3.5 9
Udio-130 12
Udio-32 30
Half Fake Suno-v2 51
Suno-v3 117
Suno-v3.5 67
Mostly Fake Suno-v2 258
Suno-v3 494
Suno-v3.5 196
Udio-130 286
Udio-32 700
Real YouTube 2167
Total 4440
Song Duration: The distribution of the song duration across training and test sets is shown on Fig. 6. From the figure, it can be observed that aside from a few minor outliers, the distribution of fake and real song durations across both training and test sets are similar.
Figure 6: Duration distribution between Training and Testing sets
Embedding Space: The t-SNE plots presented in Fig. 7 illustrate the data distributions within the embedding space, generated using an EfficientNetB0 [27] encoder with Mel Spectrogram audio inputs.
@@
  • Subfigure (d) depicts the distribution of fake types. Although Half Fake (Orange) songs form discernible clusters, Fully Fake (Red) songs exhibit a more scattered distribution, indicating a lack of common features. Conversely, Mostly Fake (Green) songs are spread broadly across the embedding space, reflecting their diverse characteristics.
@@
Figure 8: Data distributions of song-style between train and test split in the word cloud
Figure 9: Song Arena on Huggingface
Figure 10: Song Arena Leaderboard on Huggingface
Appendix B Implementation Details
To provide a clear understanding of the proposed SpecTTTra model, we present below the PyTorch-like code for its core component, the Spectro-Temporal Tokenizer ($st\_tokenizer$). It is important to note that although the provided code is in a functional format for clarity, an object-oriented approach was used in the actual implementation.
@@ Appendix B Implementation Details
def st_tokenizer(x, t_clip, f_clip, embed_dim): B, F, T = x.size() # Temporal tokens t_tokens = tokenizer( x, F, embed_dim, t_clip, T // t_clip, ) # Spectral tokens f_tokens = tokenizer( x.transpose(1, 2), T, embed_dim, f_clip, F // f_clip, ) # Spectro-Temporal tokens st_tokens = torch.cat( (t_tokens, f_tokens), dim=1 ) return st_tokens
def tokenizer( x, input_dim, token_dim, clip_size, n_clips ): # Slicing and Tokenization conv1d = nn.Conv1d( in_channels=input_dim, out_channels=token_dim, kernel_size=clip_size, stride=clip_size, bias=False, ) x = conv1d(x).gelu().transpose(1, 2) # Positional Embedding pos_embeds = nn.Parameter( torch.randn(1, n_clips, token_dim) * 0.02 ) x = x + pos_embeds # Layer Normalization x = nn.LayerNorm(token_dim, eps=1e-6)(x) return x

### v4

@@ 4.3.2 Human-AI Benchmark
To evaluate Human performance in comparison to AI-based models, we selected a subset of 520 samples from our large test data. This evaluation employed a dynamic scoring system, similar to LMSYS (Chiang et al., 2024), allowing public participation and live leaderboard updates, which will be made publicly available after decision of this paper. Three human participants were involved in this benchmark, with their performance summarized in Table 6. In contrast to the AI benchmark using short (5 sec) or long (120 sec) audio samples, this human benchmark employed 25 sec clips. This choice stems from the observation that short clips hinder human identification due to subtle inaudible artifacts easily detected by AI, while longer clips do not necessarily improve human performance due to how difficult it is to notice long-range temporal dependencies.
As shown in Table 6, AI-based methods consistently outperform human participants across all test partitions. However, both humans and AI models struggle most with Udio algorithms, particularly Udio 32, where human sensitivity dropped to 23%. Conversely, Suno algorithms, especially Suno v3.5, are easier to detect, with a human sensitivity of 82%. This mirrors the findings in the AI benchmark, where models demonstrated higher specificity than sensitivity, indicating greater ease in identifying real songs compared to fake ones. Further analysis revealed distinct patterns within real and fake songs. For instance, Suno algorithms often produced synthetic or mechanical-sounding vocals, while Udio 32 algorithm occasionally created the “Karaoke effect.” Furthermore, Udio algorithms demonstrated the ability to create songs with multiple voices and higher notes, a feature absent in Suno algorithms. On the other hand, real songs exhibit unique features such as a wide note range, diverse timbre, complex rhythms, clear vocals, and unique sounds like flutes and finger snaps.
Table 5: Comparison of conventional models and SpecTTTra against human evaluators.
@@
Algorithm (Sens.) Suno v2 0.65 0.80 0.65 0.54 0.64 0.72 0.69
Suno v3 0.99 0.96 0.96 0.98 0.98 0.99 0.75
Suno v3.5 0.99 0.98 0.98 0.99 0.99 0.99 0.82
Udio 32 0.67 0.56 0.58 0.18 0.23 0.33 0.23
Udio 130 0.99 0.87 0.98 0.99 0.99 0.99 0.55
Fake Type (Sens.) Half Fake 0.91 0.93 0.90 0.88 0.90 0.92 0.71
Mostly Fake 0.84 0.79 0.81 0.64 0.69 0.72 0.66
Full Fake 0.83 0.77 0.78 0.64 0.68 0.72 0.63
Real (Spec.) 0.98 0.80 0.98 0.97 0.95 0.94 0.78
Fake (Sens.) 0.85 0.82 0.82 0.66 0.72 0.75 0.66
Overall (F1) 0.92 0.82 0.87 0.78 0.80 0.83 0.71
Table 6: Comparison of SpecTTTra against conventional models on efficiency related metrics.
@@
(M) $\downarrow$
5 ConvNeXt 137 1.5 0.4 4 28
ViT 156 1.1 0.2 2 17
EfficientViT 55 0.6 0.5 5 22
SpecTTTra-$\gamma$ 154 0.7 0.1 2 17
SpecTTTra-$\beta$ 152 1.1 0.2 2 17
SpecTTTra-$\alpha$ 148 2.9 0.5 6 17
120 ConvNeXt 39 43.1 11.7 129 28
ViT 34 31.7 5.3 67 17
EfficientViT 43 15.9 14.8 138 22
SpecTTTra-$\gamma$ 97 10.1 1.6 20 24
SpecTTTra-$\beta$ 80 14.0 2.3 29 21
SpecTTTra-$\alpha$ 47 23.7 3.9 50 19
4.3.3 Efficiency Benchmark
To comprehensively evaluate the efficiency of the proposed SpecTTTra model alongside other methods, we measure various metrics across different song lengths using a P100 16GB GPU. The metrics considered include Speed (A/S $\rightarrow$ Audio per Second), Floating Point Operations (FLOPs), GPU Memory Consumption (Mem.) during the forward pass with a batch size of 12, activation count (# Act.), and parameter count (# Param.). The results are summarized in Table 6. Our analysis reveals that while ViT is the fastest model for 5 sec songs, it becomes the slowest for 120 sec songs (SpecTTTra-$\alpha$ is 38% faster) and exhibits significant memory consumption (SpecTTTra-$\alpha$ uses 26% less memory), rendering it less practical for longer sequences. However, ViT remains the most efficient in terms of parameter count across both short and long songs. On the other hand, ConvNeXt, despite its strong detection performance, becomes very resource-intensive for longer sequences. It consumes a large amount of memory and has the highest FLOPs and parameter count in that category. EfficientViT shows decent performance but with a surprisingly slow speed for short songs, which is over 2x slower than other models. However, in long songs, it shows better speed and lesser FLOPs than ViT and ConvNeXt but has the largest memory requirement and activation count. In contrast, the SpecTTTra model variants excel in their efficiency without compromising competitive performance in longer sequences. For example, in 120 sec songs, the SpecTTTra-$\gamma$ variant emerges as the fastest and most memory-efficient model, being nearly 3x faster and computationally more economical than ViT while showing competitive performance to it. Similarly, the SpecTTTra-$\beta$ variant is more than 2x faster than ViT and uses 2x less memory, all while achieving 3% higher performance. Performance increase culminates in the SpecTTTra-$\alpha$ variant, which outperforms ConvNeXt and EfficinetViT by 1% and 2% respectively, and achieves the highest F1 score of 97%. It achieves this by being 20% and 9% faster while using nearly 67% and 74% less memory, respectively. Therefore, SpecTTTra has the overall best performance while also being the most efficient model in the detection benchmark.
4.4 Ablation Study
We conduct an ablation study to highlight the importance of both temporal and spectral tokens, with the findings summarized in Table 7. Additionally, we vary the number of spectral tokens independently of temporal tokens to evaluate their impact on performance. Specifically, we change the clip size ($t$, $f$) relative to our best-performing model, SpecTTTra-$\alpha$ ($t=3,f=1$), to derive further insights. Notably, while it is possible to classify real and fake songs using only temporal tokens ($F/f=0$) or only spectral tokens ($T/t=0$), the combination of both clearly yields the best performance, underscoring their complementary nature. Furthermore, increasing the song duration consistently enhance performance for both spectral and temporal tokens, reinforcing our assertion about the significance of long-context information.
Table 7: Ablation analysis of temporal and spectral tokens on model performance.
5 sec 120 sec
Temp. Clip
Size (t)
# Temp.
Tok. (T/t)
Spec. Clip
Size (f)
# Spec.
Tok. (F/f)
F1
Temp. Clip
Size (t)
# Temp.
Tok. (T/t)
Spec. Clip
Size (f)
# Spec.
Tok. (F/f)
F1
3 128 - 0 0.76 3 1248 - 0 0.91
3 128 5 25 0.78 3 1248 5 25 0.94
3 128 3 42 0.79 3 1248 3 42 0.96
3 128 1 128 0.80 3 1248 1 128 0.97
- 0 1 128 0.75 - 0 1 128 0.92
@@ 5 Conclusion
In this paper, we introduced SONICS, a comprehensive dataset for end-to-end synthetic song detection, addressing limitations in existing datasets, such as lack of music diversity, short duration, and most importantly, the absence of end-to-end AI-generated songs. Moreover, we proposed the SpecTTTra model, which efficiently captures long-range temporal relationships in songs, achieving comparable performance to existing popular models while reducing computational costs significantly. Through extensive experiments, we established both AI-based and human benchmarks, demonstrating the dataset’s effectiveness in advancing synthetic song detection research. This work paves the way for future research to more effectively distinguish AI-generated music, thereby aiding in the preservation of human musical artistry.
6 Ethics Statement
The dataset incorporates copyrighted song data from YouTube. To comply with legal standards, we provide only YouTube links to the original songs and will publicly release only the AI-generated songs. Given that the generative models used to generate these fake songs may have been trained on copyrighted songs, there could be potential concerns regarding the copyright status of our dataset. However, even if the generative models were trained using copyrighted data, our dataset falls under the fair use policy U.S. Code (2023) according to research criteria. The same is true for additional metadata that was used to generate the dataset, including lyrics, style, etc. Thus, the use of these models and relevant metadata for generating our dataset is justified. Furthermore, the practice of using generative models (which likely have been trained on copyrighted data) to create large-scale datasets has been documented in the literature and published in peer-reviewed, widely accepted conferences. Notable examples include LLaVA-Instruct-158K (from GPT-4) Liu et al. (2024), Gpt4tools (from ChatGPT) Yang et al. (2024), Camel (from GPT-3.5 Turbo) Li et al. (2023), the Baize Dataset (from ChatGPT) Xu et al. (2023), and the JourneyBench Dataset (which uses GPT-4V and GPT-4O) Wang et al. (2024), among others.
Additionally, there might be concerns that these generated models regurgitate copyrighted music. To appease these concerns and verify whether the generated songs are identical to existing real songs, we compared all the generated songs with real songs in a pairwise manner using the cosine similarity metric with EfficientNetB0 embeddings as representations. Among the top 50 songs with the highest similarity to real songs, we manually inspected each one and found no exact matches. All these songs exhibited variation in elements such as music style, vocals, instrumentation, or other features, including the “half fake” subset, where only the lyrics were shared. Given this variation, our work also falls under fair use policies U.S. Code (2023). Finally, as these fake songs are generated through paid subscriptions that allow for the use and sharing of content, our dataset will be made publicly available under a CC BY-NC 4.0 license.
We also acknowledge issues related to bias and fairness. The dataset is currently limited to English-language songs, which affects its global applicability. Future work will address this by expanding the dataset to include more languages. Notably, a gender bias is evident, with male singers dominating the song styles, a trend that may stem from either the real songs or the Gemini 1.5 Flash model that was used to extract song styles (Half Fake and Mostly Fake songs), or GPT-4 that was used to generate song styles (Full Fake Songs). Addressing this gender bias is beyond the scope of our study, and we leave it to the community to tackle in future research.
7 Reproducibility Statement
To ensure reproducibility, we have made extensive efforts to document and share all necessary details. First, we provide the complete dataset generation process, including the end-to-end pipelines. The Appendix offers additional information, such as dataset statistics, to help better understand the data. Second, the pseudo-code for the Spectro-Temporal Tokenizer of the SpecTTTra model is presented in the Appendix. All hyperparameters, training setups, and augmentation methods are detailed in the “Implementation Details” section of both the main paper and the Appendix. Third, we include all assumptions and configurations for the benchmarks, which are available in both the main paper and the Appendix. Finally, the source code is provided in the supplementary materials, which contains detailed configurations for training, model parameters, and profiling.
@@ 7 Reproducibility Statement
Contents
Appendix A Dataset Characteristics
@@
Figure 5: Long Context Correlation. The red highlighted regions in the spectrogram and raw audio indicate repetition of the same verse ("Near, far, wherever you are, I believe that the heart does go on"), rhythms, and music. Such consistency is a characteristic of real songs and can be challenging for synthetic generation methods to replicate.
Figure 6: Duration distribution between Training and Testing sets.
Song Duration: The distribution of the song duration across training and test sets is shown on Fig. 6. From the figure, it can be observed that aside from a few minor outliers, the distribution of fake and real song duration across both training and test sets are similar.
Genre Distribution: Fig. 7 reveals that, despite slight imbalances, both real and fake songs are represented across all genres. This observation also indicates that the detectors are not overfitting to specific song genres. In other words, the models are not relying on genre information to determine whether a song is real or fake. Because if the detectors were attempting to perform genre classification as a shortcut for detecting fake songs, their performance would have deteriorated. However, the strong performance of all methods demonstrates that our models are effectively learning features beyond genre classification. This finding provides confidence that the detection capability is robust and not dependent on genre-specific cues.
Figure 7: Genre Distribution between Real and Fake Songs.
Embedding Space: The t-SNE plots presented in Fig. 8 illustrate the data distributions within the embedding space, generated using an EfficientNetB0 (Tan & Le, 2019) encoder with Mel Spectrogram audio inputs.
@@
  • Subfigure (d) depicts the distribution of fake types. Although Half Fake (Orange) songs form discernible clusters, Full Fake (Red) songs exhibit a more scattered distribution, indicating a lack of common features. Conversely, Mostly Fake (Green) songs are spread broadly across the embedding space, reflecting their diverse characteristics.
@@
Figure 9: Data distributions of song-style between train and test split in the word cloud.
Appendix B Dataset Quality Evaluation
Given the large scale of the datasets, exhaustive manual verification is not feasible. However, we have rigorously evaluated our pipeline and found that any potential errors have minimal impact on the dataset’s overall quality. For instance, to assess PyAnnotate’s performance in detecting non-vocal segments, we manually analyzed 50 random samples from our dataset. The results showed that the tool is highly accurate in distinguishing between vocals and non-vocals. While some errors may exist, their impact is negligible. Specifically, human evaluators reported no instances of songs containing only background music without vocals (i.e., no false negatives). Furthermore, the total number of samples classified as non-vocal (false positives) by PyAnnotate is only 3% of the dataset. Even if we conservatively assume that all of these are false positives—though this is not the case based on our evaluation—the overall impact on the dataset remains minimal. Additionally, to assess the accuracy of the extracted song styles, we manually analyzed 50 randomly selected real songs and compared their AI-generated styles (genre, instruments, mood, and vocal type) with their actual styles, as retrieved from the internet. Our analysis showed that the generated styles were highly accurate for genre, instruments, and mood. However, occasional inaccuracies were observed in the vocal type classification, which will not impact the primary goal of our work—fake song detection.
Appendix C Implementation Details
Dataset Cost Analysis: Generating the SONICS dataset costs \$1,055, allocated across GPT-4o (\$405), Gemini 1.5 Flash (\$80), Suno (\$390), and Udio (\$180).
SingFake Training: To demonstrate the limitations of SVDD (Singing Voice Deepfake Detection) models in detecting end-to-end fake songs, we first trained the models on the SingFake dataset, which consists of 597 training songs and 480 testing/validation songs. We then evaluated the performance of these models on our SONICS-test dataset, with the results summarized in Table 2. It is important to note that due to some inactive links in the SingFake dataset, we were able to download only 1077 songs, as opposed to the 1305 songs originally reported in the paper. For training, we maintained the same model configuration as in other experiments, with one exception: we resized the $128\times 128$ (5 sec) Mel spectrograms to $224\times 224$ to leverage the pretrained ImageNet weights, compensating for the smaller size of the SingFake dataset.
Augmentation: To enhance the robustness of our models, we apply MixUp (Zhang, 2017) augmentation with $\alpha=2.5$ and a 50% probability. Additionally, we utilize SpecAugment (Park et al., 2019), applying two time masks of size 8 and one frequency mask of size 8, each with a 50% probability.
Benchmark: For the efficiency benchmark, we utilized a single P100 GPU for all experiments. To measure the inference time of each model, we performed 5 warm-up runs followed by 100 test runs with a batch size of 1 to record the processing time. The results were averaged and then inverted to determine the inference speed of each model. To compute GPU memory consumption, we used a batch size of 14 across all models, measuring the peak memory usage during a single forward pass. For calculating FLOPs, we employed the fvcore (FAIR, 2023) library.
Model: For the proposed SpecTTTra model, we use the vit_small_patch16 configuration from timm (Wightman, 2019) library with an embedding dimension of 384, 6 attention heads, 12 Transformer layers, and an MLP ratio of 2.67.
To clarify the core components of the SpecTTTra model, we provide PyTorch-like code for the Spectro-Temporal Tokenizer ($st_{t}okenizer$) below. Note that while the code is presented in a functional format for clarity, the actual implementation follows an object-oriented approach.
@@ Appendix C Implementation Details
def st_tokenizer(x, t_clip, f_clip, embed_dim): B, F, T = x.size()
# Temporal tokens t_tokens = tokenizer(x, F, embed_dim, t_clip, T // t_clip)
# Spectral tokens f_tokens = tokenizer( x.transpose(1, 2), T, embed_dim, f_clip, F // f_clip, )
# Spectro-Temporal tokens st_tokens = torch.cat((t_tokens, f_tokens), dim=1) return st_tokens
def tokenizer(x, input_dim, token_dim, clip_size, n_clips): # Slicing and Tokenization conv1d = nn.Conv1d( in_channels=input_dim, out_channels=token_dim, kernel_size=clip_size, stride=clip_size, bias=False, ) x = conv1d(x).gelu().transpose(1, 2)
# Positional Embedding pos_embeds = nn.Parameter(torch.randn(1, n_clips, token_dim) * 0.02) x = x + pos_embeds
# Layer Normalization x = nn.LayerNorm(token_dim, eps=1e-6)(x) return x